# Ringdown signatures of Kerr black holes immersed in a magnetic field

### v1

@@ Abstract
We analyze the quasinormal mode spectrum for Kerr black holes surrounded by an asymptotically uniform magnetic field, modeled with the Ernst-Wild geometry. A perturbative expansion in both the rotation parameter $a$ and the magnetic field $B$ allows separation of the perturbation equations, and we obtain the spectrum for a variety of scalar quasinormal modes over a range of parameters using the continued fraction method. We then interpolate the low-lying mode spectrum to construct an Ernst-Wild template for the ringdown, and use the LIGO-Virgo-KAGRA analysis tool pyRing to assess the impact of the magnetosphere on the extraction of ringdown signatures from several observed binary black hole mergers.
@@ 1 Introduction
A distinct, but related, question is how the basic Kerr template for these modes could be impacted theoretically by the black hole environment. Early examples include scalar hair [22, 23], and black hole charge within the full Kerr-Newman geometry [24, 25]; see also [26, 27]. While these generalizations were considered primarily as a way to test for any deviation from the Kerr spectrum, in this paper we will explore the impact of a more physical black hole environment, namely an asymptotically uniform magnetosphere as an approximation to the field structure sustained by an accretion disk. While an astrophysically relevant magnetosphere is not expected to produce large corrections to the Kerr QNM spectrum, and thus seems impractical to detect for black holes in the LVK window given the current level of precision [28], it may still provide a nontrivial ‘nuisance’ parameter that could impact the extraction of mass and spin within the Kerr waveform template. In particular, the QNM pseudospectrum is known to be highly (and nonlinearly) sensitive to perturbations of the underlying potential [29, 30, 21]. Accordingly, our main goal will be to explore the impact of a magnetosphere on the modeling of ringdown from a merger via quasinormal modes.
@@
Black hole magnetospheres have been analyzed extensively, with perhaps the most notable motivation being their potential role as power sources for active galaxies, e.g. via the Blandford-Znajek (BZ) mechanism [31]. However, such energy extraction phenomena rely on the specifics of force-free electrodynamics in the black hole ergosphere, for which no analytic solutions are known beyond the perturbative regime of slow rotation. Thus, in considering ringdown, it is of interest to study cases where exact solutions do exist, and where the back-reaction on the black hole geometry can be fully taken into account. While the magnetized Kerr-Newman solution is technically the simplest, a more physical system requires an approximately uniform field away from the black hole. The regime where the energy density in the magnetic field is dominant is described by the Melvin solution [32], in which the spacetime asymptotics are also modified, which in general will completely alter the behaviour of the QNM spectrum. A Kerr black hole immersed in an asymptotically uniform magnetic field is described by the Ernst-Wild solution [33, 34], generalizing the static Ernst solution [35]. The QNM spectrum for this system was studied in [36] focusing on the regime of very large magnetic fields, which provide box-like boundary conditions leading to the possibility of superradiance [37, 38, 39]. Our focus will instead be on the regime of small magnetic fields (relative to the black hole mass), which is more readily realized astrophysically.
A full perturbative analysis of the Ernst-Wild geometry is complicated by the fact that there is no known decoupling of the mode equations, which therefore need to be analyzed via complex numerical techniques [4, 40, 41] even in the scalar case. Since our focus concerns instead the impact on the extraction of ringdown parameters from binary mergers, in this exploratory study, we will make a number of simplifying assumptions. Firstly, we will analyze scalar perturbations, which while unphysical are known in the Kerr limit to behave similarly to the tensor modes. This correspondence is discussed further in Section 2. Secondly, we will work to leading order in both the black hole rotation parameter, and the magnitude of the magnetic field. The latter assumption is physically reasonable for astrophysical black holes, but the former assumption is instead made for technical reasons, as it allows for complete decoupling of the perturbation equations. We quantify the impact of this assumption in application to mergers in the LVK catalogue with rotation parameters up to $a/M\sim 0.7$ in the next section. One further benefit is that the mode asymptotics as $r\rightarrow\infty$ reduce to those of Kerr, rather than Melvin, so that the modes are primarily sensitive to the localized impact of the magnetic field near the Kerr horizon and photon sphere.
@@
Within the approximation scheme outlined above, we determine the scalar Ernst-Wild QNM spectrum as a function of the rotation parameter $a$ and the magnetic field $B^{2}$ for the low-lying modes most relevant for the ringdown signature. We then construct an Ernst-Wild ringdown template and utilize the public LVK ringdown pipeline pyRing to assess its fit to a number of merger events in the catalogue. More precisely, we vary the magnetic field and explore its impact on extracting the primary parameters of mass and spin for the final black hole remnant. We conclude that while the effect is not large, it warrants consideration in a similar manner to the potential impact of nonlinear corrections to the ringdown.
@@
The rest of this paper is structured as follows. Section 2 reviews the Ernst-Wild geometry for the Kerr black hole immersed in an asymptotically uniform magnetic field, and the equation for scalar quasinormal mode perturbations. Section 3 determines several quasinormal mode frequencies to leading order in the rotation and magnetic field parameters, and provides an interpolation that is then used to extrapolate over an extended range of remnant parameters. Section 4 describes the ensuing construction of an EW template, its implementation within the pyRing pipeline, and explores how a magnetosphere impacts the extraction of remnant mass and spin from several binary merger events in the LVK catalogue. The results are briefly discussed in Section 5, along with an outlook for future improvements in the analysis.
@@ 2 Perturbations of the Ernst-Wild solution
where $B$ is the magnitude of the asymptotically uniform magnetic field strength. We will assume that the magnetic field surrounding the black hole is ‘weak’, namely that $BM\ll 1$, a condition that is readily satisfied by known astrophysical magnetospheres. On restoring units, this limit takes the form [49, 50, 51],
@@ 2.1 Linearized perturbation analysis
Substitution of this ansatz into Eq. 8, and using the orthogonality of the spherical harmonics, allows the reduction to a system of nine coupled radial equations [36]. We further assume that the magnetic field surrounding the black hole is weak, $\tilde{B}=BM\ll 1$, and truncating the system of coupled radial equations at ${\cal O}(\tilde{a}^{2},\tilde{B}^{2})$, we have
$\frac{d^{2}\bar{R}_{\ell m}}{dr_{*}^{2}}=\sum_{i=-1}^{1}V_{\ell+2i,m}\bar{R}_{\ell+2i,m},$ (10)
@@
    Perturbative analysis at ${\cal}(\tilde{a}\tilde{B}^{2})$: A perturbative expansion in $\tilde{B}$ is well-motivated physically on the grounds that astrophysical magnetospheres are not expected to approach $BM{\sim}1$. Moreover, the large $B$ asymptotics are those of the Melvin universe which exhibits normal (as opposed to quasinormal) modes, significantly impacting the spectrum analyzed in [36]. A perturbative expansion in $\tilde{a}$ is less well-motivated physically, as the LVK catalogue centers on remnant black holes with $\tilde{a}\sim 0.7{-}0.8$. However, when combined with a perturbative expansion in $\tilde{B}$, this approximation allows for separable perturbations, and numerically even for $\tilde{a}\sim 0.7{-}0.8$, the QNM spectrum for the linearized geometry agree quantitatively to $\sim$25with the exact Kerr modes. This comparison is illustrated in Fig. 2, and explored further in [53]. The fractional errors are larger for the imaginary part than the real part of the modes. To formalize the expansion, we can introduce a small dimensionless parameter $\epsilon$, scaling $\tilde{a}\rightarrow\tilde{a}\epsilon^{2}$ and $\tilde{B}^{2}\rightarrow\tilde{B}^{2}\epsilon$, and then truncate the expansion at ${\cal O}(\epsilon^{3})$ which retains terms of ${\cal O}(\tilde{B}^{2},\tilde{a},\tilde{a}\tilde{B}^{2})$. Terms of ${\cal O}(\tilde{B}^{4})$ and ${\cal O}(\tilde{B}^{6})$, while parametrically included under this scaling regime, do not qualitatively change the geometry at ${\cal O}(\tilde{a})$ and are expected to be very small astrophysically, so will be neglected below.
Figure 2: The real (left) and imaginary (right) parts of the $(\ell,m,n)=\{(2,2,0),(2,2,1),(3,2,0),(4,2,0)\}$ scalar quasinormal mode frequencies for the Kerr solution as a function of $\tilde{a}$, compared to the corresponding modes of the linearized Ernst-Wild geometry (for $B\rightarrow 0$).
Figure 3: The real (left) and imaginary (right) parts of the $(\ell,m,n)=\{(2,2,0),(2,2,1),(3,2,0),(4,2,0)\}$ scalar quasinormal mode frequencies for the Ernst solution as a function of $\tilde{B}$, compared to the corresponding modes of the linearized Ernst-Wild geometry (for $\tilde{a}\rightarrow 0$).
@@ 3 QNM spectrum via continued fractions
$R_{\ell m}(r)\sim\begin{dcases}(r-2M)^{\sqrt{2M}\sqrt{\tilde{a}m\omega-2M\Omega^{2}-\rho}},&\mathrm{\ as\ }r\rightarrow 2M,\\ r^{\frac{\rho+i\left(4M\Omega^{2}\right)}{2\Omega}}e^{i\Omega r},&\mathrm{\ as\ }r\rightarrow\infty,\end{dcases}$ (21)
@@
$R_{\ell m}(r)$ $=r^{\frac{\rho+i\left(4M\Omega^{2}\right)}{2\Omega}}(r-2M)^{-\sqrt{2M}\sqrt{\tilde{a}m\omega-2M\Omega^{2}-\rho}}$
@@
Figure 4: Quasinormal mode frequencies of the linearized Ernst-Wild geometry are presented for $0\leq\tilde{a},\tilde{B}\leq 0.3$ in increments of 0.01, exhibiting the $\omega_{\ell-mn}=-\omega^{*}_{\ell mn}$ pairing for $\ell,m=\pm 1,\pm 2$ that results from axial symmetry. Starting from the bottom of the plots, as $\tilde{a}$ increases the lines move in opposite directions horizontally while as $\tilde{B}$ increases the values become less damped and move vertically upwards. These modes, including both positive and negative branches, were computed with the continued fraction methods described in Section 3
@@ 3.2 Numerical QNM interpolation
where $\omega_{\ell 0n}$ is the Schwarzchild QNM for a given  and $\ell\text{ and }n$ as listed in Table 1. We compute the coefficients $a_{\ell mn}$, $b_{\ell mn}$, and $c_{\ell mn}$ by taking numerical derivatives of the data obtained in the preceding section. In particular, $a_{\ell mn}$ describe the first derivative of the Kerr spectrum in the linearized regime of $\tilde{a}$, $b_{\ell mn}$ describe the first derivative of the Ernst spectrum in the linearized regime of $\tilde{B}^{2}$, and $c_{\ell mn}$ result from the numerical second derivative with respect to both $\tilde{a}$ and $\tilde{B}^{2}$. The results for these coefficients are tabulated in Tables 2 and 3.
@@ 4.1 Constructing the ringdown template
Here, $\omega_{s{\ell mn}}$ are the complex tensor ($s=-2$) ringdown frequencies of the Kerr black hole determined by the dimensionless final spin $\tilde{a}_{f}$ and the detector frame final mass given by $M_{f}$ defined in [75]. $M_{f}$ is related to the source frame mass by a factor of $(1+z)$, and $D_{L}$ is the luminosity distance. The $f$ subscript on parameters in this section indicates that these values refer to the final black hole of the merger, commonly referred to as the remnant.
@@
4.2 Time-domain likelihood
[…]
4.3 Ringdown analysis and discussion

### v2

@@ Abstract
We analyze the quasinormal mode spectrum for Kerr black holes surrounded by an asymptotically uniform magnetic field, modeled with the Ernst-Wild geometry. A perturbative expansion in both the rotation parameter $a$ and the magnetic field $B$ allows the analysis of perturbations with Kerr-like asymptotics well inside the Melvin radius, and we obtain the spectrum for a variety of scalar quasinormal modes over a range of parameters using the continued fraction method. We then interpolate the low-lying mode spectrum to construct an Ernst-Wild template for the ringdown, and use the LIGO-Virgo-KAGRA analysis tool pyRing to assess the impact of the magnetosphere on the extraction of ringdown signatures from several observed binary black hole mergers.
@@ 1 Introduction
A distinct, but related, question is how the basic Kerr template for these modes could be impacted theoretically by the black hole environment. Early examples include scalar hair [22, 23], and black hole charge within the full Kerr-Newman geometry [24, 25]; see also [26, 27, 28]. While these generalizations were considered primarily as a way to test for any deviation from the Kerr spectrum, in this paper we will explore the impact of a more physical black hole environment, namely an asymptotically uniform magnetosphere as an approximation to the field structure sustained by an accretion disk. While an astrophysically relevant magnetosphere is not expected to produce large corrections to the Kerr QNM spectrum, and thus seems impractical to detect for black holes in the LVK window given the current level of precision [29], it may still provide a nontrivial ‘nuisance’ parameter that could impact the extraction of mass and spin within the Kerr waveform template. In particular, the QNM spectrum is known to be highly (and nonlinearly) sensitive to perturbations of the underlying potential, as characterized by the pseudospectrum for example [30, 31, 21]. Accordingly, our main goal will be to explore the impact of a magnetosphere on the modeling of ringdown from a merger via quasinormal modes.
@@
Black hole magnetospheres have been analyzed extensively, with perhaps the most notable motivation being their potential role as power sources for active galaxies, e.g. via the Blandford-Znajek (BZ) mechanism [32]. However, such energy extraction phenomena rely on the specifics of force-free electrodynamics in the black hole ergosphere, for which no analytic solutions are known beyond the perturbative regime of slow rotation. Thus, in considering ringdown, it is of interest to study cases where exact solutions do exist, and where the back-reaction on the black hole geometry can be fully taken into account. While the magnetized Kerr-Newman solution is technically the simplest, a more physical system requires an approximately uniform field away from the black hole. The regime where the energy density in the magnetic field is dominant is described by the Melvin solution [33], in which the spacetime asymptotics are also modified, which in general will completely alter the behaviour of the QNM spectrum. A Kerr black hole immersed in an asymptotically uniform magnetic field is described by the Ernst-Wild solution [34, 35], generalizing the static Ernst solution [36]. The QNM spectrum for this system was studied in [37] focusing on the regime of very large magnetic fields, where the Melvin-like asymptotics at large radius lead to the possibility of superradiance [38, 39, 40]. Our focus will instead be on the regime of small magnetic fields (relative to the black hole mass), which is more readily realized astrophysically.
To make contact with astrophysical systems, where the magnetosphere is localized and the spacetime is asymptotically flat, we choose to expand perturbatively in the magnetic field strength $B$. On one hand, this is consistent with the small scale of observed astrophysical magnetic fields (in Planck units), but it also creates an important scale separation in the geometry. The non-asymptotically flat Melvin spacetime only becomes dominant outside the Melvin radius $r_{M}\sim 1/B\gg r_{+}$. Thus, for parametrically small $B$, the solution exhibits conventional Kerr-like asymptotics within a large intermediate interval in radius $r_{+}\ll r\ll r_{B}$. More technically, the equations for perturbations of the spacetime exhibit a transition from Kerr-like asymptotics at ${\cal O}(B^{2})$ to Melvin-like confining asymptotics at ${\cal O}(B^{4})$. As a result, at ${\cal O}(B^{2})$ the modes are primarily sensitive to the localized impact of the magnetic field near the Kerr horizon and photon sphere. Analyzing the spectrum of perturbations at this order also has the secondary benefit of technical simplicity. The analysis of linearized perturbations in the Ernst-Wild geometry is complicated by the fact that there is no known decoupling of the mode equations, which therefore need to be analyzed via complex numerical techniques [4, 41, 42, 43] even in the scalar case. However, working at ${\cal O}(B^{2})$ leads to simplifications, which we exploit in this exploratory study by also working perturbatively in the Kerr rotation parameter $a$. This allows for complete decoupling of the perturbation equations.
@@
We proceed to determine the scalar Ernst-Wild QNM spectrum as a function of the rotation parameter $a$ and the magnetic field $B^{2}$ for the low-lying modes most relevant for the ringdown signature. For simplicity, we choose to focus on scalar perturbations which, while unphysical, are known in the Kerr limit to behave similarly to the tensor modes. We quantify the impact of this assumption and the perturbative expansion in $B^{2}$ and $a$ in the next section, focusing on the application to mergers in the LVK catalogue. Using the QNM spectrum, we then construct an Ernst-Wild ringdown template and utilize the public LVK ringdown pipeline pyRing to assess its fit to a number of merger events in the catalogue. More precisely, we vary the magnetic field and explore its impact on extracting the primary parameters of mass and spin for the final black hole remnant. We conclude that while the effect is not large, it warrants consideration in a similar manner to the potential impact of nonlinear corrections to the ringdown.
@@
The rest of this paper is structured as follows. Section 2 reviews the Ernst-Wild geometry for the Kerr black hole immersed in an asymptotically uniform magnetic field, and the equation for scalar quasinormal mode perturbations. Further details of the large radius asymptotics within the perturbative expansion in $B$ are presented in Appendix A. Section 3 determines several quasinormal mode frequencies to leading order in the rotation and magnetic field parameters, and provides an interpolation that is then used to extrapolate over an extended range of remnant parameters. Section 4 describes the ensuing construction of an EW template, its implementation within the pyRing pipeline, and explores how a magnetosphere impacts the extraction of remnant mass and spin from several binary merger events in the LVK catalogue. The results are briefly discussed in Section 5, along with an outlook for future improvements in the analysis.
@@ 2 Perturbations of the Ernst-Wild solution
where $B$ is the magnitude of the asymptotically uniform magnetic field strength. We will assume that the magnetic field surrounding the black hole is ‘weak’, namely that $BM\ll 1$, a condition that is readily satisfied by known astrophysical magnetospheres. As discussed above, this restriction also has the benefit of restoring Kerr-like asymptotics for perturbations when we formally truncate the geometry at ${\cal O}(\tilde{B}^{2})$, as illustrated in Appendix A. On restoring units, this limit takes the form [51, 52, 53],
@@ 2.1 Linearized perturbation analysis
Substitution of this ansatz into Eq. 8, and using the orthogonality of the spherical harmonics, allows the reduction to a system of nine coupled radial equations [37]. As noted above, we further assume that the magnetic field surrounding the black hole is weak, $\tilde{B}=BM\ll 1$, and truncating the system of coupled radial equations at ${\cal O}(\tilde{a}^{2},\tilde{B}^{2})$, we have
$\frac{d^{2}\bar{R}_{\ell m}}{dr_{*}^{2}}+\sum_{i=-1}^{1}V_{\ell+2i,m}\bar{R}_{\ell+2i,m}=0,$ (10)
@@
    Perturbative analysis at ${\cal}(\tilde{a}\tilde{B}^{2})$: A perturbative expansion in $\tilde{B}$ is well-motivated physically on the grounds that astrophysical magnetospheres are not expected to approach $BM{\sim}1$. Moreover, the large $B$ asymptotics are those of the Melvin universe which exhibits normal (as opposed to quasinormal) modes, significantly impacting the spectrum analyzed in [37]. As described in Appendix A, the restriction to ${\cal O}(\tilde{B}^{2})$ restores the Kerr-like asymptotic boundary conditions, and ensures that the QNM only reflect the impact of the magnetic field near the black hole. A perturbative expansion in $\tilde{a}$ is less well-motivated physically, as the LVK catalogue centers on remnant black holes with $\tilde{a}\sim 0.7{-}0.8$. However, when combined with a perturbative expansion in $\tilde{B}$, this approximation allows for separable perturbations, and numerically even for $\tilde{a}\sim 0.7{-}0.8$, the QNM spectrum for the linearized geometry agrees quantitatively to $\sim$25with the exact Kerr modes. This comparison is illustrated in Fig. 2 (see also [54, 55]). The fractional errors are larger for the imaginary part than the real part of the modes. To formalize the expansion, we can introduce a small dimensionless parameter $\epsilon$, scaling $\tilde{a}\rightarrow\tilde{a}\epsilon^{2}$ and $\tilde{B}^{2}\rightarrow\tilde{B}^{2}\epsilon$, and then truncate the expansion at ${\cal O}(\epsilon^{3})$ which retains terms of ${\cal O}(\tilde{B}^{2},\tilde{a},\tilde{a}\tilde{B}^{2})$. Terms of ${\cal O}(\tilde{B}^{4})$ and ${\cal O}(\tilde{B}^{6})$, while parametrically allowed under this scaling regime, are small inside the Melvin radius for the $\tilde{B}$-range considered here and are excluded to ensure Kerr-like large radius asymptotics as described above.
Figure 2: The real (left) and imaginary (right) parts of the $(\ell,m,n)=\{(2,2,0),(2,2,1)\}$ scalar quasinormal mode frequencies for the Kerr solution as a function of $\tilde{a}$, compared to the corresponding modes of the linearized Ernst-Wild geometry (for $B\rightarrow 0$). The differences for the $\{(3,2,0),(4,2,0)\}$ modes are very similar.
@@ 3 QNM spectrum via continued fractions
$R_{\ell m}(r)\sim\begin{dcases}(r-2M)^{-\sqrt{2M}\sqrt{\tilde{a}m\omega-2M\Omega^{2}-\rho}},&\mathrm{\ as\ }r\rightarrow 2M,\\ r^{\frac{i\left(\rho+4M\Omega^{2}\right)}{2\Omega}}e^{i\Omega r},&\mathrm{\ as\ }r\rightarrow\infty,\end{dcases}$ (21)
@@
$R_{\ell m}(r)$ $=r^{\frac{i\left(\rho+4M\Omega^{2}\right)}{2\Omega}}r^{\sqrt{2M}\sqrt{\tilde{a}m\omega-2M\Omega^{2}-\rho}}$
×$\times(r-2M)^{-\sqrt{2M}\sqrt{\tilde{a}m\omega-2M\Omega^{2}-\rho}}$
@@
Figure 3: Quasinormal mode frequencies of the linearized Ernst-Wild geometry are presented for $0\leq\tilde{a},\tilde{B}\leq 0.3$ in increments of 0.01, exhibiting the $\omega_{\ell-mn}=-\omega^{*}_{\ell mn}$ pairing for $\ell,m=\pm 1,\pm 2$ that results from axial symmetry. Starting from the bottom of the plots, as $\tilde{a}$ increases the lines move in opposite directions horizontally while as $\tilde{B}$ increases the values become less damped and move vertically upwards. These modes, including both positive and negative branches, were computed with the continued fraction methods described in Section 3.
@@ 3.2 Numerical QNM interpolation
Figure 4: The real (left) and imaginary (right) parts of the $(\ell,m,n)=\{(2,2,0),(2,2,1)\}$ scalar quasinormal mode frequencies for the Ernst solution as a function of $\tilde{B}$, compared to the corresponding modes within the linearized fit to Ernst-Wild spectrum in (35) (for $\tilde{a}\rightarrow 0$). The differences for the $\{(3,2,0),(4,2,0)\}$ modes are very similar.
@@
where $\omega_{\ell 0n}$ is the Schwarzchild QNM for a given  and $\ell\text{ and }n$ as listed in Table 1. We compute the coefficients $a_{\ell mn}$, $b_{\ell mn}$, and $c_{\ell mn}$ by taking numerical derivatives of the data obtained in the preceding section. In particular, $a_{\ell mn}$ describe the first derivative of the Kerr spectrum in the linearized regime of $\tilde{a}$, $b_{\ell mn}$ describe the first derivative of the Ernst spectrum in the linearized regime of $\tilde{B}^{2}$, and $c_{\ell mn}$ result from the numerical second derivative with respect to both $\tilde{a}$ and $\tilde{B}^{2}$. The results for these coefficients are tabulated in Tables 2 and 3, with those restricted to linearized Kerr agreeing with [54]. A comparison between the modes obtained directly from the Ernst geometry at ${\cal O}(\tilde{B}^{2})$ and the coefficients $b_{lmn}$ is shown in Fig. 4.
@@ 4.1 Constructing the ringdown template
Here, $\tilde{\omega}_{s{\ell mn}}$ are the complex tensor ($s=-2$) ringdown frequencies of the Kerr black hole determined by the dimensionless final spin $\tilde{a}_{f}$ and the detector frame final mass given by $M_{f}$ defined in [77]. $M_{f}$ is related to the source frame mass by a factor of $(1+z)$, and $D_{L}$ is the luminosity distance. The $f$ subscript on parameters in this section indicates that these values refer to the final black hole of the merger, commonly referred to as the remnant.
@@
4.2 Time-domain likelihood
[…]
4.3 Ringdown analysis and discussion
@@ Acknowledgements.
Appendix A Ernst-Wild asymptotics at ${\cal O}(\tilde{B}^{2})$
In this Appendix, we provide further details of the large-$r$ asymptotic behaviour of Ernst-Wild perturbations when working perturbatively at a fixed order in $\tilde{B}$. For this purpose, it is helpful to first consider the pure Melvin spacetime, which determines the large-$r$ asymptotics of the Ernst-Wild geometry. The line element is given by setting $M=\tilde{a}=0$ in Eq. 2,
$ds^{2}_{M}=\Lambda\left(-dt^{2}+dr^{2}+r^{2}d\theta^{2}\right)+\frac{\sin^{2}\theta}{\Lambda}d\phi^{2},$ (A41)
where $\Lambda=\left(1-\frac{1}{4}B^{2}r^{2}\sin^{2}\theta\right)^{2}$. In the absence of the black hole, it is more convenient to present the Melvin line element using cylindrical coordinates $\rho=r\sin\theta$ and $z=r\cos\theta$,
$ds^{2}_{M}=\Lambda(-dt^{2}+d\rho^{2}+dz^{2})+\frac{\rho^{2}}{\Lambda}d\phi^{2},$ (A42)
with $\Lambda=\left(1-\frac{1}{4}B^{2}\rho^{2}\right)^{2}$. This solution describes the back-reaction to a uniform magnetic field aligned along the $z$-direction, and reflects the non-asymptotically flat nature of the Ernst-Wild geometry at large radius. Following the approach of Section 2, we can solve for massless scalar perturbations using Eq. 8 with the following cylindrical ansatz,
$\Phi(t,\rho,z,\phi)=\frac{R(\rho)}{\sqrt{\rho}}e^{-i\omega t}e^{im\phi}e^{ikz},$ (A43)
and the Klein Gordon equation becomes [37],
$\begin{split}R^{\prime\prime}(\rho)&-\frac{m^{2}\left(B^{2}\rho^{2}+4\right)^{4}R(\rho)}{256\rho^{2}}\\ &-\frac{64\left(4k^{2}\rho^{2}-4\rho^{2}\omega^{2}-1\right)R(\rho)}{256\rho^{2}}=0.\end{split}$ (A44)
The behaviour at infinity for the exact Melvin spacetime is given by
$R(\rho\rightarrow\infty)\rightarrow(B\rho)^{-3/2}e^{-(B\rho)^{4}m/64},$ (A45)
reflecting the confining asymptotics. However, in the regime $B\rho\ll 1$, and neglecting terms at $\mathcal{O}(B^{4})$, the asymptotic boundary condition instead reduces to the form
$R(\rho\rightarrow\infty)|_{{\cal O}(B^{2})}\rightarrow e^{i\rho\sqrt{\omega^{2}-k^{2}-B^{2}m^{2}}},$ (A46)
equivalent to an outgoing wave in Minkowski space, in contrast to the full Melvin asymptotics. Thus, provided we work perturbatively at ${\cal O}(B^{2})$, perturbations are largely insensitive to the non-asymptotically flat geometry that sets in at large distance, and more precisely beyond the Melvin radius $\rho_{M}=2/B$. It follows that, after adding a black hole at the origin in this coordinate system, only the localized changes to the geometry near the horizon modify the mode solutions at ${\cal O}(B^{2})$.
Returning to the full Ernst-Wild solution, the perturbative analysis above reflects the nontrivial transition in asymptotics in going from ${\cal O}(\tilde{B}^{2})$ to ${\cal O}(\tilde{B}^{4})$. Due to the long-range $1/r$ potential, the asymptotic boundary condition (21) for perturbations at ${\cal O}(\tilde{B}^{2})$ is equivalent to that for the Kerr spacetime, with the only correction appearing as a phase shift. Within the approximation (A2) outlined in Section 2, we have neglected terms of ${\cal O}(\tilde{B}^{4})$ that appear in the form $\tilde{B}^{4}r^{4}$, thus the perturbation equation given by Eq. 18 in the main text applies in the intermediate regime
$r\ll r_{M}\sim 1/B.$ (A47)
We therefore have a notion of ‘infinity’ close to $r/M=1/\tilde{B}$, meaning that the model implicitly matches the Ernst-Wild metric far from the black hole to an asymptotically flat solution. However, when considering the perturbations that describe quasinormal modes this formal matching procedure is not necessary because the equation already exhibits the Kerr asymptotics due to the restriction to ${\cal O}(\tilde{B}^{2})$.
Figure 7: Radial effective potential for $(\ell,m)=(1,1)$ for the Schwarzschild metric ($\tilde{B}=0$) and the Ernst metric ($\tilde{B}=0.05$). The shaded region is beyond the Melvin radius $\sim 1/B$, although the potential only receives a constant shift as $r\rightarrow\infty$ when restricted to ${\cal O}(\tilde{B}^{2})$.
Concretely, we can compare the perturbation potentials of the static (and non-asymptotically flat) Ernst geometry (with $\tilde{a}=0$) and the Schwarzschild geometry (with $\tilde{a}=\tilde{B}=0$) via Eq. 19 in the main text,
$\begin{split}V_{\ell m}|_{{\cal O}(\tilde{B}^{2})}=&-\frac{(2M-r)\left(\ell(\ell+1)r+2M\right)}{\,r^{4}}\\ &-\frac{(2M-r)\tilde{B}^{2}m^{2}r^{3}}{M^{2}r^{4}}.\end{split}$ (A48)
The effective potential for the Schwarzschild black hole on the first line is only corrected by a constant at large radius at ${\cal O}(\tilde{B}^{2})$. The shift is exhibited in Fig. 7, with the shaded region showing the domain beyond the Melvin radius where the asymptotics would be modified more substantially at ${\cal O}(\tilde{B}^{4})$. The restriction to ${\cal O}(\tilde{B}^{2})$ thus restores the expectation within the Kerr spacetime that the QNM spectrum is primarily affected by the behaviour of the potential near the horizon and photon sphere of the black hole [98, 99, 63, 100, 37].